OpenStudy (anonymous):

1. Take any number (except for 1). Square that number and then subtract one. Divide by one less than your original number. Now subtract your original number. Did you reach 1 for an answer? You should have. 2. How does this number game work? Hint. Redo the number game using a variable instead of an actual number and rewrite the problem as one rational expression. The number I picked was 28, how do I rewrite this as a rational expression?

OpenStudy (anonymous):

\[(x^2-1)/(x-1)-x=(x+1)(x-1)/(x-1)-x=(x+1)-x=1\]

OpenStudy (anonymous):

you rewrite it as a rational expression with a variable (I used x)

OpenStudy (anonymous):

The goal of a rational expression is to equal one?

OpenStudy (anonymous):

If you solve the expression that is described in words, calling x the initial number, you get 1 as a solution for any value of x (except 1, since if x=1, x-1=0 and you can't divide by 0)

OpenStudy (anonymous):

The rational expression is the first thing I wrote (before the first =) (x2−1)/(x−1)−x

OpenStudy (anonymous):

I guess I am confused on what a rational expression is ?

OpenStudy (anonymous):

I would call that an expression with ratios (i.e. something divided by something else)

OpenStudy (anonymous):

But beware english is NOT my mother language!

OpenStudy (anonymous):

Okay, I am still a lil confused, but I will keep studying.Thanks!